OpenStudy (anonymous):

Goofy Golf ticket prices for kids under 12 years of age increased from $4.00 to $6.00 this year. What is the percent of increase in the tickets? 50% 60% 40%

Parth (parthkohli):

Percentage increase = (Increase/Original) * 100

Parth (parthkohli):

= (2/4) * 100 = 0.5 * 100 can you do it now?

OpenStudy (anonymous):

50%

Parth (parthkohli):

yes

OpenStudy (anonymous):

thx